# Nonlinear modulation of dispersive fast magnetosonic waves in an inhomogeneous rotating solar low-$\beta$ magnetoplasma

### v1

@@ III Derivation of NLS equation
where $v_{g}$ is the group velocity of the wave envelope along the $x$-axis and $\epsilon$ is a small ($0<\epsilon\ll 1$) expansion parameter, which scales the weakness of amplitudes of perturbations. Due to the stretched coordinates (17), the space and time derivatives will be replaced according to the following transformations.
@@ III.1 First-order perturbations: Linear dispersion relation
Figure 2 displays the profiles of the real wave mode ($\Re\omega$, for $k>k_{J}$) and the growth rate of instability ($\Im\omega$, for $k<k_{J}$) by the effects of the density and magnetic field inhomogeneities [See Eq. (15)], the Coriolis force, and the self-gravity force such that $\omega_{J}>2\Omega_{0}$. We find that while the real wave frequency increases with the wave number $k$, the instability growth rate falls off from a nonzero value with $k$ having a cut-off at the critical Jeans wave number, i.e., $k=k_{J}$. Such a critical value shifts towards a higher value of $k$ due to the effects of the inhomogeneities in which $\psi_{0}(x)$ increases, but $\rho_{0}(x)$ decreases with increasing values of $x>0$. In this case, the instability domain for $k$ expands and the domain of the real wave mode reduces. Thus, it follows that the density and magnetic field inhomogenities in the background plasma favor the Jeans instability in a wide range of values of $k$, not reported before. On the other hand, a small reduction of the angular frequency $\Omega_{0}$ of the rotating fluid, associated with the Coriolis force, can significantly reduce both the growth rate of instability and the instability domain in $k$. We do not consider the case of $\omega_{J}\approx 2\Omega_{0}$ at which the fast magnetosonic wave becomes dispersionless. Such a strict condition may be applicable to low-frequency long-wavelength plasma oscillations.
@@ IV MODULATIONAL INSTABILITY
where $\eta_{0}$ is a constant, and $\eta_{j}$ and $\sigma_{j}$ for $j=1,2$, are the amplitudes of perturbations.
@@
In Subsections IV.1-IV.3, we will study the conditions for the modulational instability and the instability growth rate in the three different cases as mentioned above.
@@ IV.2 Case II: $\omega_{J}>2\Omega_{0}$, $k>k_{J}$
Figure 4 displays the growth rate of instability [Eq. (36)] for different parameter values as in Fig. 2. We find that the instability growth rate can be enhanced and maximized by the effects of the magnetic field or density inhomogeneity. However, it can be reduced or minimized by reducing the angular frequency $\Omega_{0}$ compared to the Jeans frequency.
@@ IV.3 Case III: $\omega_{J}>2\Omega_{0}$, $k<k_{J}$
Figure 6: The instability growth rates (a) $\Gamma_{1}$ [Eq. (40), when $P_{1}Q_{1}<0$] $\Gamma_{2}$ [Eq. (42), when $P_{1}Q_{1}>0$] are plotted against the modulation wave number $K$ of perturbation. The fixed parameter values and different parameter values the solid, dashed, and dotted lines are the same as for Fig. 2.
@@ V Applications and discussions
Typical solar plasma parameters are [35] (i) the number density of electrons/ions, $n\sim 10^{9}~{}\rm{cm}^{-3}$, (ii) the electron temperature, $T_{e}\sim 10^{-3}$ K, and (iii) the plasma $\beta\sim 0.2~{}(\beta\rightarrow 0)$ at the magnetic field strength, $B_{0}\sim 4$ G  ($40$ G). For illustration purpose, we calculate the maximum growth rate in the case of $\omega_{J}>2\Omega_{0}$, $k>k_{J}$ (Case II in Sec. IV.2) at $B_{0}=4$ G as $\Gamma_{\rm{max}}\sim 0.0076$ (for $\eta_{0}=0.01$) such that the typical growth time can be estimated as $\tau_{\rm{growth}}\sim 4$ s, i.e., faster than predicted before [27] and may be reasonable to explain the transition time (twenty minutes) from fast magnetosonic mode to slowly varying envelopes. The typical Jeans wavelength and the Alfvén velocity can be estimated as $\lambda_{J}\sim 2\times 10^{4}$ m and $V_{A}\sim 3\times 10^{5}$ m/s. We can also estimate the group velocity as $v_{g}\sim 5\times 10^{5}$ m/s. These velocities are close to the observational velocity $190$ km/s of slowly moving perturbations. The instability growth rate can be significantly high by the effects of the magnetic field and density inhomogeneities where the Alfvén velocity $V_{A}(x)$ gets reduced. Note that the above results are valid for low-$\beta$ plasmas where the length scale of magnetic field inhomogeneity is much smaller than that of the density inhomogeneity and the wavelength associated with the fast carrier magnetosonic mode is small compared to the wavelength of the slow modulation along the magnetic field. On the other hand, the typical growth rate can be high having no cut-off at any wavelength of modulation, and the corresponding time scale can be low in the case when $\omega_{J}>2\Omega_{0}$ and $k<k_{J}$ (Case III in Sec. IV.2). However, to be consistent with the small perturbations, the admissible growth rate should be at a wavelength (of slow modulation) above its critical value.
@@
It is to be noted that in the case of plasmas with constant gravity, the gravity force, $\mathbf{g}=-\nabla\psi$ is to be replaced by $\mathbf{g}=(-g,0,0)$ and there will be no gravitational Poisson equation. Even if it is considered for the sake of clarity, it will result into the quasineutrality condition, already assumed for the derivation of the MHD equations. In this case, the term $-\omega_{J}^{2}$ in the dispersion equation (22) for the fast magnetosonic carrier waves will be replaced by $igk$, which will result into the real wave frequency, given by,
@@ VI Summary and conclusion
To summarize, we have studied the dispersion properties of fast magnetosonic waves and the modulation of slowly varying magnetosonic wave envelopes in inhomogeneous magnetoplasmas under the relative influence of the gravitational force and the Coriolis force due to rotating fluids. Using the multiple-scale reductive perturbation technique, we have derived the NLS equation for the evolution of slowly varying magnetosonic envelopes. It is found that the inhomogeneities in the background magnetic field and the fluid density favor the Jeans instability to occur in a larger domain of the carrier wave number $(k)$ below the Jeans critical wave number $(k_{J})$. The latter is, however, defined when the self-gravity force dominates over the Coriolis force or the Jean frequency $\omega_{J}$ becomes larger than the increased angular frequency $2\Omega_{0}$. On the other hand, when the Coriolis force dominates over the self-gravity force or is dominated by the self-gravity force with $k>k_{J}$, the fast magnetosonic carrier mode is always stable. It is interesting to note that the MSW envelope, corresponding to the unstable carrier Jeans mode with $\omega_{J}>2\Omega_{0}$ and $k<k_{J}$, is always unstable under the plane wave modulation having no cut-offs in the growth rates. However, the stable Jeans mode with $\omega_{J}>2\Omega_{0}$ but $k>k_{J}$ can lead to modulational stability or instability with finite growth rates and cut-offs. Such an instability growth rate is enhanced by the effects of the magnetic field or density inhomogeneities. We have also discussed the possible applications of modulational instability in solar plasmas, such as those in the X-ray corona and solar prominence, and noted that the estimated Alfvén velocity of fast modes and the group velocity of slowly varying envelopes are close to the observational values. Also, the estimated growth time is found to be reasonable to explain the transition time from the fast magnetosonic modes to slowly varying envelopes at the nonlinear stage.
@@ Acknowledgments
One of us, J. Turi, wishes to thank the Council of Scientific and Industrial Research (CSIR) for a Senior Research Fellowship (SRF) with reference number 09/202(0115)/2020-EMR-I. The authors also thank Sima Roy of the University of Engineering & Management, Kolkata, for rechecking the derivations of different expressions for the NLS equation.

### v2

@@ II Basic equations
It is to be mentioned that in a general manner, Eqs. (1)-(4) describe three independent propagating modes, namely, the Alfvén wave and the fast and slow magnetosonic waves. These waves can be distinguished by their speeds and polarizations. In the present work, we have chosen the propagation direction along the x-axis, i.e., perpendicular to the external magnetic field $B_{0}\hat{z}$. Thus, we have only the fast magnetosonic wave, since the slow magnetosonic wave and the Alfvén wave become non-propagating and degenerate. However, if we consider the polarization in a plane and propagation direction along an axis or as arbitrary having components along the axes, there would appear both the Alfvén and magnetosonic waves as fundamental wave modes, which may be coupled, e.g., due to magnetic field line curvature [32], the Hall effects [33] and/or the Coriolis force for rotating fluids [31]. The nonlinear interactions of these high-frequency Alfvén waves and low-frequency magnetosonic waves can give rise to several interesting phenomena, including wave structures [32] and MHD wave turbulence due to energy exchange [34].
@@ III Derivation of NLS equation
where $v_{g}$ is the group velocity of the wave envelope along the $x$-axis and $\epsilon$ is a small ($0<\epsilon\ll 1$) expansion parameter, which scales the weakness of amplitudes of perturbations. Due to the stretched coordinates (17), the space and time derivatives will be replaced according to the following transformations:
@@
We use the reductive perturbation technique to derive the NLS equation (with second-order group-velocity dispersion and cubic or Kerr nonlinearity) for the nonlinear evolution of first-order first harmonic (for $n=1,~{}l=1$) mode or the fundamental mode, i.e., the fast magnetosonic wave. In the linear limit, $l=1$ gives the dispersion relation for the $(\omega,k)$ mode with $n=1$ and the group velocity (second-order) dispersion with $n=2$. However, the second-order zeroth $(l=0)$ harmonic modes can appear due to nonlinear self-interactions of the fundamental modes with complex conjugates ($l=1$ and $l=-1$). In the NLS equation, the cubic nonlinearity for the fundamental mode, i.e., of the form $|A_{1}^{(1)}|^{2}A_{1}^{(1)}$, appears in the nonlinear interactions between the (i) zeroth harmonic second-order modes $(l=0,~{}n=2)$ and first harmonic first-order modes or fundamental modes $(l=1,~{}n=1)$ and (ii) second harmonic second-order modes $(l=2,~{}n=2)$ and the complex conjugate of the first harmonic first-order modes or fundamental modes $(l=-1,~{}n=1)$. Thus, the second harmonic second order modes $(l=2,~{}n=2)$ while interacting with the complex conjugate of the fundamental mode can generate the cubic nonlinearity in the evolution equation for the fundamental mode. Although the analysis for the derivation of the NLS equation is standard and well-known, for clarity, we have added a diagram to show routes from linear to nonlinear couplings of different modes (See Fig. 2).
Figure 2: A schematic diagram showing routes to obtain the linear dispersion relation and the group velocity dispersion of the fundamental mode (first-order first harmonic) and the cubic nonlinearity due to nonlinear couplings of the fundamental modes.
@@ III.1 First-order perturbations: Linear dispersion relation
Figure 3 displays the profiles of the real wave mode ($\Re\omega$, for $k>k_{J}$) and the growth rate of instability ($\Im\omega$, for $k<k_{J}$) by the effects of the density and magnetic field inhomogeneities [See Eq. (15)], the Coriolis force, and the self-gravity force such that $\omega_{J}>2\Omega_{0}$. We find that while the real wave frequency increases with the wave number $k$, the instability growth rate falls off from a nonzero value with $k$ having a cut-off at the critical Jeans wave number, i.e., $k=k_{J}$. Such a critical value shifts towards a higher value of $k$ due to the effects of the inhomogeneities in which $\psi_{0}(x)$ increases, but $\rho_{0}(x)$ decreases with increasing values of $x>0$. In this case, the instability domain for $k$ expands and the domain of the real wave mode reduces. Thus, it follows that the density and magnetic field inhomogenities in the background plasma favor the Jeans instability in a wide range of values of $k$, not reported before. On the other hand, a small reduction of the angular frequency $\Omega_{0}$ of the rotating fluid, associated with the Coriolis force and the Jeans frequency $\omega_{J}$, can significantly reduce both the growth rate of instability and the instability domain in $k$. We do not consider the case of $\omega_{J}\approx 2\Omega_{0}$ at which the fast magnetosonic wave becomes dispersionless. Such a strict condition may be applicable to low-frequency long-wavelength plasma oscillations.
@@ IV MODULATIONAL INSTABILITY
where $\eta_{0}$ is a constant, and $\eta_{j}$ and $\sigma_{j}$ for $j=1,2$, are the amplitudes of perturbations. Thus, the solution (31) after the modulation according to Eq. (34) can be expressed as
$\begin{split}B(x,t)=&\frac{1}{2}\sqrt{\eta_{0}}\cos(k_{0}x-\omega_{0}t)\\ &+\chi_{1}(K)\cos\left[\left(k_{0}+\epsilon K\right)x-\left(\omega_{0}+\epsilon v_{g}K+\epsilon^{2}\Omega\right)t\right]\\ &+\chi_{2}(K)\sin\left[\left(k_{0}+\epsilon K\right)x-\left(\omega_{0}+\epsilon v_{g}K+\epsilon^{2}\Omega\right)t\right]\\ &+\rm{similar~{}terms~{}with}~{}(-K,-\Omega),\end{split}$ (35)
where $\chi_{j}(K),~{}j=1,2$ is given by
$\chi_{j}(K)=\frac{1}{4}\left(\frac{i}{\sqrt{\eta_{0}}}+\frac{\sqrt{\eta_{0}}}{PK}\sigma_{j}\right).$ (36)
From Eq. (35), it is evident that the modulation of magnetosonic wave envelopes is basically a process of three-wave interaction among the unperturbed carrier wave $(k_{0},\omega_{0})$ and the two side bands with the wave numbers $k_{0}\pm\epsilon K$ and wave frequencies $\omega_{0}\pm\epsilon v_{g}K\pm\epsilon^{2}\Omega$. Accordingly, the modulational instability (MI) in a nonlinear medium is also known as sideband instability. In general, such instabilities can result from different kinds of nonlinearities in the media. Similar to nonlinear fiber optics, where the MI occurs due to anomalous chromatic dispersion and the Kerr nonlinearity of an optical fiber, the MI in dispersive plasma media is caused by the cubic or Kerr nonlinearity $(Q)$ of an electrostatic or electromagnetic pulse in conjunction with the frequency-dependent group velocity dispersion $(P)$ (i.e., no higher-order dispersion, no propagation loss). Consequently, the sidebands get amplified in the wave spectrum, leading to increasing oscillations of the wave amplitude. The physical origin of MI of fast magnetosonic waves is the magnetic field-driven ponderomotive force, which can act as a negative pressure to accumulate plasmas in the region where the fast magnetosonic modes exist.
@@
From the dispersion relation [Eq. (37)] it can be remarked that the frequency becomes imaginary and hence the perturbation grows exponentially (instability) or the small sinusoidal amplitude modulation can be amplified if $PQ>0$ and $K<K_{c}\equiv\sqrt{2\eta_{0}Q/P}$ (in the case when $P$ and $Q$ are reals). The gain spectrum can be defined by $\Gamma\equiv\Im\Omega$ so that the power of a wave perturbation grows with time as $\sim\sqrt{\eta}\exp(\Gamma\tau)$. In Subsections IV.1-IV.3, we will study the conditions for the modulational instability and the instability growth rate in the three different cases as mentioned above.
@@ IV.2 Case II: $\omega_{J}>2\Omega_{0}$, $k>k_{J}$
Figure 5 displays the growth rate of instability [Eq. (38)] for different parameter values as in Fig. 3. We find that the instability growth rate can be enhanced and maximized by the effects of the magnetic field or density inhomogeneity. However, it can be reduced or minimized by reducing the angular frequency $\Omega_{0}$ and the Jeans frequency.
@@ IV.3 Case III: $\omega_{J}>2\Omega_{0}$, $k<k_{J}$
Figure 7: The instability growth rates (a) $\Gamma_{1}$ [Eq. (42), when $P_{1}Q_{1}<0$] and (b) $\Gamma_{2}$ [Eq. (44), when $P_{1}Q_{1}>0$] are plotted against the modulation wave number $K$ of perturbation. The fixed parameter values and different parameter values the solid, dashed, and dotted lines are the same as for Fig. 3.
@@ V Applications and discussions
Typical solar coronal plasma parameters are [38] (i) the number density of electrons/ions, $n\sim 10^{9}~{}\rm{cm}^{-3}$, (ii) the electron temperature, $T_{e}\sim 10^{6}$ K, and (iii) the plasma $\beta\sim 0.2~{}(\beta\rightarrow 0)$ at the magnetic field strength, $B_{0}\sim 4$ G  ($40$ G). For illustration purpose, we calculate the maximum growth rate in the case of $\omega_{J}>2\Omega_{0}$, $k>k_{J}$ (Case II in Sec. IV.2) at $B_{0}=4$ G as $\Gamma_{\rm{max}}\sim 0.0076$ (for $\eta_{0}=0.01$) such that the typical growth time can be estimated as $\tau_{\rm{growth}}\sim 4$ s, i.e., faster than predicted before [27] and may be reasonable to explain the transition time (twenty minutes) from fast magnetosonic mode to slowly varying envelopes. The typical Jeans wavelength and the Alfvén velocity can be estimated as $\lambda_{J}\sim 2\times 10^{4}$ m and $V_{A}\sim 3\times 10^{5}$ m/s. We can also estimate the group velocity as $v_{g}\sim 5\times 10^{5}$ m/s. These velocities are close to the observational velocity $190$ km/s of slowly moving perturbations. The instability growth rate can be significantly high by the effects of the magnetic field and density inhomogeneities where the Alfvén velocity $V_{A}(x)$ gets reduced. Note that the above results are valid for low-$\beta$ plasmas where the length scale of magnetic field inhomogeneity is much smaller than that of the density inhomogeneity and the wavelength associated with the fast carrier magnetosonic mode is small compared to the wavelength of the slow modulation along the magnetic field. On the other hand, the typical growth rate can be high having no cut-off at any wavelength of modulation, and the corresponding time scale can be low in the case when $\omega_{J}>2\Omega_{0}$ and $k<k_{J}$ (Case III in Sec. IV.2). However, to be consistent with the small perturbations, the admissible growth rate should be at a wavelength (of slow modulation) above its critical value.
@@
While the self-gravity force works, e.g., in the inner layers of the Sun: The core, the radiative zone, and the convective zone, or the cores of giant stars, the constant gravity may apply to regions of planetary atmosphere (e.g., Earth’s atmosphere) or solar atmosphere [e.g., Sun’s atmosphere (outer): Photosphere, Chromosphere, and Corona]. Thus, in the case of plasmas with constant gravity, the gravity force, $\mathbf{g}=-\nabla\psi$ is to be replaced by $\mathbf{g}=(-g,0,0)$ and there will be no gravitational Poisson equation. Even if it is considered for the sake of clarity, it will result into the quasineutrality condition, already assumed for the derivation of the MHD equations. In this case, the term $-\omega_{J}^{2}$ in the dispersion equation (22) for the fast magnetosonic carrier waves will be replaced by $igk$, which will result into the real wave frequency, given by,
@@ VI Summary and conclusion
To summarize, we have studied the dispersion properties of fast magnetosonic waves and the modulation of slowly varying magnetosonic wave envelopes in inhomogeneous magnetoplasmas under the relative influence of the gravitational force and the Coriolis force due to rotating fluids. Using the multiple-scale reductive perturbation technique, we have derived the NLS equation for the evolution of slowly varying magnetosonic envelopes. It is found that the inhomogeneities in the background magnetic field and the fluid density favor the Jeans instability to occur in a larger domain of the carrier wave number $(k)$ below the Jeans critical wave number $(k_{J})$. The latter is, however, defined when the self-gravity force dominates over the Coriolis force or the Jean frequency $\omega_{J}$ becomes larger than the increased angular frequency $2\Omega_{0}$. On the other hand, when the Coriolis force dominates over the self-gravity force or is dominated by the self-gravity force with $k>k_{J}$, the fast magnetosonic carrier mode is always stable. It is interesting to note that the MSW envelope, corresponding to the unstable carrier Jeans mode with $\omega_{J}>2\Omega_{0}$ and $k<k_{J}$, is always unstable under the plane wave modulation having no cut-offs in the growth rates. However, the stable Jeans mode with $\omega_{J}>2\Omega_{0}$ but $k>k_{J}$ can lead to modulational stability or instability with finite growth rates and cut-offs. Such an instability growth rate is enhanced by the effects of the magnetic field or density inhomogeneities. It is important to note that in self-gravitating rotating plasmas, when the Coriolis force dominates over the self-gravity force ($\omega_{J}<2\Omega_{0}$), the carrier Jeans magnetosonic mode is always stable and so is the modulation perturbation of magnetosonic envelopes. In this case, the Coriolis force plays a role in stabilizing both the linear (Jeans) and nonlinear (modulational) instabilities. On the other hand, when the Coriolis force is dominated by the gravity force ($\omega_{J}>2\Omega_{0}$ and $k<k_{J}$), the Coriolis force itself or both the forces can not stabilize the instabilities but can reduce the corresponding growth rates of instabilities either at higher values of $\Omega_{0}$ or at lower values of both $\omega_{J}$ and $\Omega_{0}$. However, both the forces can stabilize the modulational instability at lower values of both $\omega_{J}$ and $\Omega_{0}$ even when $\omega_{J}>2\Omega_{0}$ but $k>k_{J}$ (the case where the carrier Jeans mode is linearly stable). We have also discussed the possible applications of modulational instability in solar plasmas, such as those in the X-ray corona and solar prominence, and noted that the estimated Alfvén velocity of fast modes and the group velocity of slowly varying envelopes are close to the observational values. Also, the estimated growth time is found to be reasonable to explain the transition time from the fast magnetosonic modes to slowly varying envelopes at the nonlinear stage.
@@
Modulational instability is a prerequisite for energy localization and the structure formation of different types of envelope solitons in the NLS equation. The unstable and stable modulation domains reported here could be helpful to look for bright and dark envelope solitons, Kuznetsov–Ma breather soliton, Akhmediev breather, as well as Peregrine solitons [41]. We have limited our discussion to the one-dimensional propagation of magnetosonic waves and the modulation along the carrier wave vector. However, considering the arbitrary direction of propagation and the magnetic field in a plane can give rise to Alfvén and magnetosonic modes as fundamental modes that can get coupled via the Coriolis force for rotating fluids [31] or due to Hall effects in Hall MHD plasmas [33]. In this case, any one (to be modified by the other) can be a fundamental mode to study the modulational instability of wave envelopes. The amplitude modulation obliquely to the wave vector could result in different instability criteria with stable and unstable regions [42] that would differ from those presented here. Also, in multi-dimensional propagation of magnetosonic waves, the modulation along a wave vector may not be described by a single NLS-type equation but a coupled set of equations, i.e., Davey–Stewartson-like equations [42]. The latter can govern the possible collapse of magnetosonic wave envelopes. Thus, oblique modulation of multi-dimensional magnetosonic wave envelopes in Hall MHD rotating plasmas could be a project for future studies.
@@ Acknowledgments
One of us, J. Turi, wishes to thank the Council of Scientific and Industrial Research (CSIR) for a Senior Research Fellowship (SRF) with reference number 09/202(0115)/2020-EMR-I. The authors also thank Sima Roy of the University of Engineering & Management, Kolkata, for rechecking the derivations of different expressions for the NLS equation and the anonymous referees for some helpful comments, which improved the manuscript in its present form.